# Categorification of generic Su-Zhang Character Formula

### v1

@@
(December 18, 2025)
For semisimple Lie algebras, the BGG resolution is commonly interpreted as a categorification of the Weyl character formula. For general linear Lie superalgebras, the Kac–Wakimoto character formula is regarded as a good character formula for a certain class of ”singular” weights (the so-called Kac–Wakimoto weights). However, the combinatorial factor $\frac{1}{\operatorname{atyp}(\lambda)!}$ makes a direct BGG-type categorification of this formula rather elusive.
In this paper we construct, for a certain class of “generic” weights, a resolution that categorifies a known Weyl-type finite-sum character formula of the same flavour as the Kac–Wakimoto formula. Our resolution is built from taking images of canonical homomorphisms between Verma modules corresponding to non-conjugate Borel subalgebras via odd reflections.
Note that, while the infinite resolution of Kac–Wakimoto simple modules by Kac modules constructed by Brundan–Stroppel [brundan2010highestII] does not directly generalize the classical BGG resolution, the construction developed here does.
@@ 1 Introduction
General linear Lie superalgebras $\mathfrak{gl}(m|n)$ are widely regarded as natural generalizations of general linear Lie algebras, and their representation theory has been extensively studied as a rich and interesting subject. One of the most fundamental problems is to determine the characters of finite-dimensional simple representations. This problem was solved independently by Serganova [serganova1996kazhdan] and Brundan [brundan2003kazhdan]. Their results can be expressed as the following infinite sum:
@@
where $L^{()}(\lambda)$ denotes the simple highest weight module, $K_{\lambda,\mu}(-1)$ is the value at 1-1 of Serganova-Brundan’s Kazhdan–Lusztig-type polynomial (now well understood in terms of the arc–diagram calculus of Brundan–Stroppel [stroppel2012highest]), and $K_{\geq 0}(\mu)$ is the Kac module, i.e. the standard object in the finite-dimensional weight module category $\mathcal{F}$.
@@
Building on Brundan’s result [brundan2003kazhdan], Su and Zhang [su2007character] obtained, for the first time, a Weyl-type character formula given by a finite sum. Unlike in the purely even Lie algebra case, however, the Su–Zhang character formula is rather intricate and not very convenient for explicit computations. We refer to [gorelik2023gruson] for a detailed overview of the history of character formulas for Lie superalgebras. Nevertheless, in several important classes one can obtain very neat closed character formulas [chmutov2014weyl]. The best-known example is the Kac–Wakimoto character formula (conjectured in [kac1994integrable] and proved in [chmutov2015kac, cheng2017kac, gorelik2015characters]). Very roughly, for a so-called Kac–Wakimoto weight (a.k.a. Kostant weight or totally disconnected weight) $\lambda$ one has
@@
$\operatorname{ch}L^{()}(\lambda)=\frac{1}{\operatorname{atyp}(\lambda)!}\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})},$
@@
where $M^{()}(w\cdot\lambda)$ is the Verma module, $\Gamma_{\lambda}$ denotes the set of atypical odd positive roots for $\lambda$ and $\operatorname{atyp}(\lambda)$ denotes the atypicality of $\lambda$.
@@
In the classical setting of semisimple Lie algebras, the Weyl character formula is categorified by the BGG resolution [bgg1973schubert]. It is thus natural to ask whether one can construct a resolution that categorifies the Kac–Wakimoto character formula. The presence of the factor $\frac{1}{\operatorname{atyp}(\lambda)!}$, however, makes such a categorification problem rather delicate. Note that Brundan–Stroppel [brundan2010highestII] constructed, generalizing [cheng2008bgg], for Kac–Wakimoto weights $\lambda$, an explicit and very interesting resolution of the Kac module $K_{\geq 0}(\lambda)$ in terms of diagram algebras. This should be viewed as a categorification of the Serganova–Brundan infinite-sum formula, rather than of a Weyl-type formula. See also [coulembier2014bernstein].
@@
Kac–Wakimoto weights can be considered, in a sense, as the most singular among regular dominant weights. On the other hand, there are natural “generic” classes of weights. In this paper we work with the class of $\mathfrak{g}_{-1}$-generic weights. Our notion of $\mathfrak{g}_{-1}$-genericity coincides with the $\Gamma^{+}$–genericity of Coulembier [coulembier2016bott], and is therefore weaker than the notion of weakly generic and generic in Penkov-Serganova [coulembier2016primitive, penkov1989cohomology, penkov1994generic]. In particular, any weight lying outside a small neighbourhood of the walls of the Weyl chambers is $\mathfrak{g}_{-1}$–generic.
@@
Su and Zhang [su2007character] showed that, on a certain generic region which contains all $\mathfrak{g}_{-1}$-generic weights, their character formula simplifies to
@@
$\operatorname{ch}L^{()}(\lambda)=\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}.$
@@
Here the troublesome factor $\frac{1}{\operatorname{atyp}(\lambda)!}$ disappears. (See also [chmutov2014weyl] for a nice exposition.) This makes it natural to try to construct a BGG-type resolution for such $\lambda$. More precisely, we aim to construct a “narrow” Verma module $N^{()}(\lambda)$ whose character is given by
$\operatorname{ch}N^{()}(\lambda)=\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})},$
and then to build a resolution of $L^{()}(\lambda)$ in terms of the modules.
@@
Our main point of view is that the above narrow Verma modules and their resolutions can be constructed from odd reflections and changes of Borel subalgebras. In contrast to the situation for semisimple Lie algebras, a basic classical Lie superalgebra admits several non-conjugate Borel subalgebras, and passing from one Borel to another via odd reflections is a genuinely new and essential feature of the theory. From the perspective of Nichols algebras of diagonal type [heckenberger2009classification, andruskiewitsch2017finite] and their classification via their Weyl groupoids (a.k.a. generalized root systems) [heckenberger2008generalization, heckenberger2020hopf, bonfert2024weyl], these changes of Borel are also of independent interest.
@@
Our narrow Verma modules are defined as follows. Let $()$ be the distinguished Borel subalgebra and $(n^{m})$ the antidistinguished Borel subalgebra of $\mathfrak{gl}(m|n)$. We define the narrow Verma module $N^{()}(\lambda)$ to be the image of up to scalar, a unique nonzero homomorphism Verma modules sharing same character with different choice of Borels;
@@
Moreover,
$\operatorname{ch}N^{()}(\lambda)=\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}.$
@@ 2 Well known basics
From now on, we will denote by $\mathfrak{g}$ a finite dimensional Lie superalgebra. We consider the category $\mathfrak{g}\text{-sMod}$, where morphisms respects $\mathbb{Z}/2\mathbb{Z}$-grading. (This is the module category of a monoid object in $\mathfrak{g}\text{-sMod}$ in the sense of [etingof2015tensor].)
@@ 2.1 Root systems $\Delta$
It is clear that the root system of $\mathfrak{gl}(m|n)$, after forgetting the parity of roots, identifies with that of $\mathfrak{gl}(m+n)$. In particular, it is useful to consider the Weyl group $S_{m+n}$ of $\mathfrak{gl}(m+n)$. This phenomenon is specific to $\mathfrak{gl}(m|n)$ and does not occur for other basic classical Lie superalgebras.
@@
The even positive root systems are all conjugate under $S_{m}\times S_{n}$. We therefore fix the standard even Borel $\mathfrak{b}^{\mathrm{st}}_{\bar{0}}$ (block upper triangular in $\mathfrak{g}_{\bar{0}}$), with even positive root system
$\Delta_{\bar{0}}^{+}=\{\ \varepsilon_{i}-\varepsilon_{j}\mid 1\leq i<j\leq m\ \}\ \cup\ \{\ \delta_{p}-\delta_{q}\mid 1\leq p<q\leq n\ \}.$
With this choice fixed, positive systems whose even part equals $\Delta_{\bar{0}}^{+}$ are classified by $\varepsilon\delta$–sequences, equivalently by $(m|n)$-shuffle
$\tau\in\mathrm{Sh}(m|n):=\bigl\{\,w\in S_{m+n}\ \big|\ w(1)<\cdots<w(m)\ \text{and}\ w(m+1)<\cdots<w(m+n)\,\bigr\}.$
Given $\tau$, set
$\Delta^{+}(\tau)=\{\ \varepsilon_{i}-\varepsilon_{j}\mid\tau(i)<\tau(j)\ \}$
yielding a bijection $\tau\leftrightarrow\Delta^{+}(\tau)$; hence there are $\frac{(m+n)!}{m!n!}$ positive systems with standard even one.
Definition 2.2.
A partition is a weakly decreasing finite sequence of positive integers; we allow the empty partition (). For example, we write $5+3+3+1$ as $(53^{2}1)$. We identify partitions with Young diagrams in French notation (rows increase downward). A box has coordinates $(i,j)$ with $i$ the column (from the left) and $j$ the row (from the bottom).
Each $\varepsilon\delta$–sequences with $n$ symbols $\varepsilon$ and $m$ symbols $\delta$ determines a lattice path from $(n,0)$ to $(0,m)$ (left step for $\varepsilon$, up step for $\delta$); the region weakly southeast of the path inside $(n^{m})$ is a Young diagram fitting in the $m\times n$ rectangle $(n^{m})$, and this gives a bijection between $\varepsilon\delta$–sequences and the set of Young diagrams fitting in the $m\times n$ rectangle $(n^{m})$.
In what follows, we represent a Borel subalgebra with standard even Borel subalgebra by a partition. In particular, we call () the distinguished Borel subalgebra, and $(n^{m})$ the anti-distinguished Borel subalgebra. We denote by $L(m,n)$ the set of all Borel subalgebras of $\mathfrak{gl}(m|n)$ whose even part is the standard Borel subalgebra of $\mathfrak{g}_{\overline{0}}$.
@@ 2.2 Contragredient duality $()^{\vee}$
Lemma 2.4 (See also [brundan2014representations] Lemma 2.2, [chen2020primitive] Proposition 2.2.3).
@@
Lemma 2.5 ([bell1993theory], Th. 2.2; [coulembier2017gorenstein], §6.1).
@@
Note that $\Lambda^{\dim\mathfrak{g}_{\overline{1}}}(\mathfrak{g}_{\overline{1}})\cong\Lambda^{2mn}(\mathfrak{g}_{\overline{1}})\cong L_{\bar{0}}(0)$ in our general linear setting, but in general, we have $\operatorname{Ind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\cong\operatorname{Coind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\circ(-\otimes\Lambda^{\dim\mathfrak{g}_{\bar{1}}}\mathfrak{g}_{\bar{1}})$. $L_{\bar{0}}(0)=\Lambda^{0}(\mathfrak{g}_{\overline{1}})$ and $\Lambda^{\dim\mathfrak{g}_{\overline{1}}}(\mathfrak{g}_{\overline{1}})$ are direct summands of $\Lambda(\mathfrak{g}_{\overline{1}})$.
@@ 2.3 $\mathfrak{b}$-Verma modules $M^{\mathfrak{b}}(\lambda)$
Let $\mathfrak{b}\in L(m,n)$. Let $k_{\lambda}^{\mathfrak{b}}$ be the one-dimensional $\mathfrak{b}$-module corresponding to $\lambda\in\Lambda$. The $\mathfrak{b}$-Verma module with highest weight $\lambda$ is defined by $M^{\mathfrak{b}}(\lambda)=U(\mathfrak{g})\otimes_{U(\mathfrak{b})}k_{\lambda}^{\mathfrak{b}}.$ (Here, the parity of $k_{\lambda}^{\mathfrak{b}}$ is chosen so that $M^{\mathfrak{b}}(\lambda)\in\mathcal{W}$. )
@@
Definition 2.8 (integral Weyl vectors [brundan2014representations]).
@@
We define the $\mathfrak{b}$-atypicality of $\lambda$ as
@@
If $\mathrm{aty}^{\mathfrak{b}}(\lambda)=0$, then $\lambda$ is called $\mathfrak{b}$-typical; otherwise, it is called $\mathfrak{b}$-atypical.
[…]
@@
Lemma 2.11.
For $\mathfrak{b},\mathfrak{b}^{\prime}\in L(m,n)$ and $\lambda,\lambda^{\prime}\in\Lambda$, the following statements hold:
  1. 1.
    $\operatorname{ch}M^{\mathfrak{b}}(\lambda-\rho^{\mathfrak{b}})=\operatorname{ch}M^{\mathfrak{b}^{\prime}}(\lambda-\rho^{\mathfrak{b}^{\prime}})$;
  2. 2.
    $\dim M^{\mathfrak{b}^{\prime}}(\lambda-\rho^{\mathfrak{b}^{\prime}})_{\lambda-\rho^{\mathfrak{b}}}=1$;
  3. 3.
    $\operatorname{ch}M^{\mathfrak{b}}(\lambda)=\operatorname{ch}M^{\mathfrak{b}^{\prime}}(\lambda^{\prime})\iff\lambda+\rho^{\mathfrak{b}}=\lambda^{\prime}+\rho^{\mathfrak{b}^{\prime}}$;
  4. 4.
    $\dim\operatorname{Hom}(M^{\mathfrak{b}}(\lambda-\rho^{\mathfrak{b}}),M^{\mathfrak{b}^{\prime}}(\lambda-\rho^{\mathfrak{b}^{\prime}}))=1$;
@@ 2.4 Kac functors $K_{\geq 0}$
Define the inflation functor $\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\geq 0}}\colon\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}_{\geq 0}\text{-sMod}$ by letting $\mathfrak{g}_{1}$ act trivially. Define the functor $(\,)^{\mathfrak{g}_{1}}\colon\mathfrak{g}_{\geq 0}\text{-sMod}\to\mathfrak{g}_{\bar{0}}\text{-sMod}$ by $M\mapsto M^{\mathfrak{g}_{1}}:=\{\,m\in M\mid\mathfrak{g}_{1}m=0\,\}$ . Then these functors are mutually adjoint each other and exact.
The Kac functor is defined as $K_{\geq 0}:=\mathrm{Ind}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}\circ\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\geq 0}}:\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}\text{-sMod}$, which is exact. Then $K_{\geq 0}$ is left adjoint to the exact functor $\mathrm{Res}_{\geq 0}:=(\,)^{\mathfrak{g}_{1}}\circ\mathrm{Res}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}:\mathfrak{g}\text{-sMod}\to\mathfrak{g}_{\bar{0}}\text{-sMod}$. If we denote by $\mathrm{Coind}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}$ [chen2021simple] the right adjoint of $\mathrm{Res}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}$, and define $\mathrm{Coind}_{\geq 0}:=\mathrm{Coind}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}\circ\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\geq 0}}:\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}\text{-sMod}$ then there is an isomorphism of functors $\mathrm{Coind}_{\geq 0}\circ(-\otimes L_{\bar{0}}(-2\rho_{\bar{1}}))\cong K_{\geq 0}$, see [chen2021simple]. Note that $L_{\bar{0}}(-2\rho_{\bar{1}})\cong\bigwedge^{\dim\mathfrak{g}_{-1}}\mathfrak{g}_{-1}$ as $\mathfrak{g}_{\bar{0}}$-modules.
@@
[germoni1998indecomposable] $K_{\geq 0}\cong(\,\cdot\,)^{\vee}\circ\mathrm{Coind}_{\leq 0}\circ(\,\cdot\,)^{\vee},\qquad K_{\leq 0}\cong(\,\cdot\,)^{\vee}\circ\mathrm{Coind}_{\geq 0}\circ(\,\cdot\,)^{\vee}.$ $K_{\geq 0}$ and $K_{\leq 0}$ restricts between $\mathcal{O}_{\bar{0}}$ and $\mathcal{O}$.
@@
( [chen2021simple]) For a weight $\lambda$, the following are true:
@@
[chen2021translated] For every $M\in\mathfrak{g}_{\bar{0}}\text{-sMod}$ which is finite length, one has $\ell(\operatorname{soc}M)=\ell(\operatorname{soc}K_{\geq 0}M)=\ell(\operatorname{soc}K_{\leq 0}M)$ and $\ell(\operatorname{top}M)=\ell(\operatorname{top}K_{\geq 0}M)=\ell(\operatorname{top}K_{\leq 0}M)$, where $\ell(-)$ denotes the length.
@@
Proposition 2.16 ([mazorchuk2014parabolic, chen2021tilting]).
@@
Theorem 2.17 ([chen2023some]).
labelthm:soc-dist Let $\lambda$ be an integral weight . Then

### v2

@@
For semisimple Lie algebras, the BGG resolution is often viewed as a categorification of the Weyl character formula. For general linear Lie superalgebras, Brundan–Stroppel [5] constructed an infinite resolution of the so-called Kostant simple modules by Kac modules, but their construction does not directly generalize the classical BGG resolution.
In this paper we construct, for weights lying outside a neighborhood of the walls of the Weyl chambers, a resolution that categorifies a known Weyl-type finite-sum character formula in the same spirit as the Kac–Wakimoto formula. Our resolution is built from images of canonical homomorphisms between Verma modules attached to non-conjugate Borel subalgebras related by odd reflections. In particular, the construction developed here does generalize the classical BGG resolution.
@@ 1 Introduction
General linear Lie superalgebras $\mathfrak{gl}(m|n)$ are widely regarded as natural generalizations of general linear Lie algebras, and their representation theory has been extensively studied as a rich and interesting subject. One of the most fundamental problems is to determine the characters of finite-dimensional simple representations. This problem was solved independently by Serganova [34] and Brundan [7]. Their results can be expressed as the following infinite sum:
@@
where $L^{()}(\lambda)$ denotes the simple highest weight module, $K_{\lambda,\mu}(-1)$ is the value at 1-1 of Serganova-Brundan’s Kazhdan–Lusztig-type polynomial (now well understood in terms of the arc–diagram calculus of Brundan–Stroppel [6]), and $K_{\geq 0}(\mu)$ is the Kac module, i.e. the standard object in the finite-dimensional weight module category $\mathcal{F}$.
@@
Building on Brundan’s result [7], Su and Zhang [36] obtained, for the first time, for any finite dimensional weight modules, a Weyl-type character formula given by a finite sum. Unlike in the purely even Lie algebra case, however, the Su–Zhang character formula is rather intricate and not very convenient for explicit computations. There are also other works on character formulas for $\mathfrak{gl}(m|n)$, such as [24], [16], and [35]. Nevertheless, in several important classes one can obtain very neat closed character formulas [16]. The best-known example is the Kac–Wakimoto character formula (conjectured in [29] and proved in [17, 15, 25]). Very roughly, for a so-called Kac–Wakimoto weight (a.k.a. Kostant weight or totally disconnected weight) $\lambda$ one has
@@
$\operatorname{ch}L^{()}(\lambda)=\frac{1}{\operatorname{atyp}(\lambda)!}\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{w\cdot\lambda}}(1+e^{-\beta})},$
@@
where $M^{()}(w\cdot\lambda)$ is the Verma module, $\Gamma_{w\cdot\lambda}$ denotes the set of atypical odd positive roots for $\lambda$ and $\operatorname{atyp}(\lambda)$ denotes the atypicality of $\lambda$.
@@
In the classical setting of semisimple Lie algebras, the Weyl character formula is categorified by the BGG resolution [3]. It is thus natural to ask whether one can construct a resolution that categorifies the Kac–Wakimoto character formula. The presence of the factor $\frac{1}{\operatorname{atyp}(\lambda)!}$, however, makes such a categorification problem rather delicate. Note that Brundan–Stroppel [5] constructed, generalizing [14], for Kac–Wakimoto weights $\lambda$, an explicit and very interesting resolution of the Kac module $K_{\geq 0}(\lambda)$ in terms of diagram algebras. This should be viewed as a categorification of the Serganova–Brundan infinite-sum formula, rather than of a Weyl-type formula. On the other hand, in the setting of Lie superalgebras, it is well known among experts that finite-dimensional simple modules cannot, in general, be resolved by Verma modules [14]. See also [19].
@@
Kac–Wakimoto weights can be considered, in a sense, as the most singular among regular dominant weights. On the other hand, there are natural “generic” classes of weights. In this paper we work with the class of $\mathfrak{g}_{-1}$-generic weights. Our notion of $\mathfrak{g}_{-1}$-genericity coincides with the $\Gamma^{+}$–genericity of Coulembier [20], and is therefore weaker than the notion of weakly generic and generic in Penkov-Serganova [18, 32, 33]. In particular, any weight lying outside a small neighbourhood of the walls of the Weyl chambers is $\mathfrak{g}_{-1}$–generic.
@@
Su and Zhang [36] showed that, on a certain generic region which contains all $\mathfrak{g}_{-1}$-generic weights, their character formula simplifies to
@@
$\operatorname{ch}L^{()}(\lambda)=\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{w\cdot\lambda}}(1+e^{-\beta})}.$
@@
Here the troublesome factor $\frac{1}{\operatorname{atyp}(\lambda)!}$ disappears. (See also [16] for a nice exposition.) This makes it natural to try to construct a BGG-type resolution for such $\lambda$.
@@
Our main point of view is that such resolutions can be constructed from odd reflections and changes of Borel subalgebras. In contrast to the situation for semisimple Lie algebras, a basic classical Lie superalgebra admits several non-conjugate Borel subalgebras, and passing from one Borel to another via odd reflections is a genuinely new and essential feature of the theory. From the perspective of Nichols algebras of diagonal type [28, 1] and their classification via their Weyl groupoids (a.k.a. generalized root systems) [27, 26, 4], these changes of Borel are also of independent interest.
@@
Our resolutionss are defined as follows. Let $()$ be the distinguished Borel subalgebra and $(n^{m})$ the antidistinguished Borel subalgebra of $\mathfrak{gl}(m|n)$. We define the narrow Verma module $N^{()}(\lambda)$ to be the image of up to scalar, a unique nonzero homomorphism Verma modules sharing same character with different choice of Borels;
@@
Moreover, for a $\mathfrak{g}_{-1}$-generic weight $\lambda$, we have,
$\operatorname{ch}N^{()}(\lambda)=\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{w\cdot\lambda}}(1+e^{-\beta})}.$
@@ 2 Well known basics
From now on, we will denote by $\mathfrak{g}$ a finite dimensional Lie superalgebra. We consider the category $\mathfrak{g}\text{-sMod}$, where morphisms respects $\mathbb{Z}/2\mathbb{Z}$-grading. (This is the module category of a monoid object in $\mathfrak{g}\text{-sMod}$ in the sense of [22].)
@@ 2.1 Root systems $\Delta$
We call the Borel subalgebra associated with the positive root system $\Delta^{()+}:=\{\varepsilon_{i}-\varepsilon_{j}\mid 1\leq i<j\leq m\}\cup\{\delta_{p}-\delta_{q}\mid 1\leq p<q\leq n\}\cup\{\varepsilon_{i}-\delta_{q}\mid 1\leq i\leq m,\ 1\leq q\leq n\}$ the distinguished Borel subalgebra, and denote it by $()$. Similarly, we call the Borel subalgebra associated with $\Delta^{(n^{m})+}:=\{\varepsilon_{i}-\varepsilon_{j}\mid 1\leq i<j\leq m\}\cup\{\delta_{p}-\delta_{q}\mid 1\leq p<q\leq n\}\cup\{\delta_{p}-\varepsilon_{i}\mid 1\leq p\leq n,\ 1\leq i\leq m\}$ the anti-distinguished Borel subalgebra, and denote it by $(n^{m})$.
@@ 2.2 Contragredient duality $()^{\vee}$
Lemma 2.3 (See also [8] Lemma 2.2, [11] Proposition 2.2.3).
@@
Lemma 2.4 ([2], Th. 2.2; [21], §6.1).
@@ 2.3 $\mathfrak{b}$-Verma modules $M^{\mathfrak{b}}(\lambda)$
Let $\mathfrak{b}$ be a Borel subalgebra. Let $k_{\lambda}^{\mathfrak{b}}$ be the one-dimensional $\mathfrak{b}$-module corresponding to $\lambda\in\Lambda$. The $\mathfrak{b}$-Verma module with highest weight $\lambda$ is defined by $M^{\mathfrak{b}}(\lambda)=U(\mathfrak{g})\otimes_{U(\mathfrak{b})}k_{\lambda}^{\mathfrak{b}}.$ (Here, the parity of $k_{\lambda}^{\mathfrak{b}}$ is chosen so that $M^{\mathfrak{b}}(\lambda)\in\mathcal{W}$. )
@@
Definition 2.7 (integral Weyl vectors [8]).
@@
We define the atypicality of a simple module $L=L^{\mathfrak{b}}(\lambda)$ as
@@
This definition is independent of the choice of $\mathfrak{b}$. If $\mathrm{aty}L=0$, then $L$ is called typical; otherwise, it is called atypical.
@@ 2.4 Kac functors $K_{\geq 0}$
Define the inflation functor $\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\geq 0}}\colon\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}_{\geq 0}\text{-sMod}$ by letting $\mathfrak{g}_{1}$ act trivially. Define the functor $(\,)^{\mathfrak{g}_{1}}\colon\mathfrak{g}_{\geq 0}\text{-sMod}\to\mathfrak{g}_{\bar{0}}\text{-sMod}$ by $M\mapsto M^{\mathfrak{g}_{1}}:=\{\,m\in M\mid\mathfrak{g}_{1}m=0\,\}$ . Then $\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\geq 0}}$ is left adjoint to $(\,)^{\mathfrak{g}_{1}}$.
The Kac functor is defined as $K_{\geq 0}:=\mathrm{Ind}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}\circ\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\geq 0}}:\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}\text{-sMod}$, which is exact. Then $K_{\geq 0}$ is left adjoint to the exact functor $\mathrm{Res}_{\geq 0}:=(\,)^{\mathfrak{g}_{1}}\circ\mathrm{Res}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}:\mathfrak{g}\text{-sMod}\to\mathfrak{g}_{\bar{0}}\text{-sMod}$. If we denote by $\mathrm{Coind}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}$ [12] the right adjoint of $\mathrm{Res}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}$, and define $\mathrm{Coind}_{\geq 0}:=\mathrm{Coind}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}\circ\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\geq 0}}:\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}\text{-sMod}$ then there is an isomorphism of functors $\mathrm{Coind}_{\geq 0}\circ(-\otimes L_{\bar{0}}(-2\rho_{\bar{1}}))\cong K_{\geq 0}$, see [12]. Note that $L_{\bar{0}}(-2\rho_{\bar{1}})\cong\bigwedge^{\dim\mathfrak{g}_{-1}}\mathfrak{g}_{-1}$ as $\mathfrak{g}_{\bar{0}}$-modules.
@@
[23] $K_{\geq 0}\cong(\,\cdot\,)^{\vee}\circ\mathrm{Coind}_{\leq 0}\circ(\,\cdot\,)^{\vee},\qquad K_{\leq 0}\cong(\,\cdot\,)^{\vee}\circ\mathrm{Coind}_{\geq 0}\circ(\,\cdot\,)^{\vee}.$
@@
( [12]) For a weight $\lambda$, the following are true:
@@
[10] For every $M\in\mathfrak{g}_{\bar{0}}\text{-sMod}$ which is finite length, one has $\ell(\operatorname{soc}M)=\ell(\operatorname{soc}K_{\geq 0}M)=\ell(\operatorname{soc}K_{\leq 0}M)$ and $\ell(\operatorname{top}M)=\ell(\operatorname{top}K_{\geq 0}M)=\ell(\operatorname{top}K_{\leq 0}M)$, where $\ell(-)$ denotes the length.
@@
Proposition 2.14 ([30, 9]).
@@
Theorem 2.15 ([13]).
Let $\lambda$ be an integral weight . Then